# Network-informed Prompt Engineering against Organized Astroturf Campaigns under Extreme Class Imbalance

### v1

@@ Abstract.
isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
To address the above issues, in this paper, we propose a novel framework for identifying coordinated disinformation campaigns relying only on frozen LLMs. The proposed framework encodes both the structural and textual information of a propagation tree as a text, ready to be prompted to an LLM. Using prompt engineering and retrieval-augmented generation (RAG) allows LLMs to detect organized astroturf campaigns on social networks such as $\mathbb{X}$. Our framework does not involve any training or fine-tuning of the given LLM. In contrast, it capitalizes on the power of prompt engineering and RAG techniques and enables LLMs to overcome the effects of class imbalance, thereby effectively identifying coordinated political campaigns. Our framework is flexible and can be coupled with any prompt engineering and RAG technique (e.g. few-shot prompting). Furthermore, it is compatible with any decoder-only LLM.
@@
Research questions. In addition, we answer the following research questions (Section 6):
@@ 3.4. Prompt engineering
Finally, a combination of prompting techniques can also be followed. We found that mixing role prompting, CoT and few-shot prompting yields the best results. A list of prompt examples considering the problem of detecting astroturf campaigns is available in Appendix D. We refer to Section 2 for a more thorough explanation of prompt engineering techniques.
@@ 3.5. Balanced Retrieval-Augmented Generation (Balanced RAG)
These balanced pairs are then incorporated into the few-shot prompting approach used to query the LLM. By providing the LLM with representative examples that are both similar to the query and equally cover both classes, we enhance the model’s ability to distinguish between positive and negative instances. This approach enables the LLM to learn the critical differences that define each class, thereby improving its ability to correctly classify new, unseen propagation trees.
@@ 4.2. Setup
Evaluation metrics. Since it is important to identify the positive samples of a dataset (i.e., fake news campaigns), we selected precision, recall, F1 score and ROC AUC scores as our main evaluation metrics. We provide further information about these metrics in Appendix C.
@@ 4.5. Hyperparameter tuning
Figure 2. Comparison of different metrics (precision, recall, F1 score, and ROC AUC) across various (Llama) model sizes. The graph illustrates how each metric changes as the number of model parameters increases. The exact numerical results are also available in Table 7.
@@ 5.2. Impact of the encoded information
6. Discussion
@@
7. Conclusion
Appendix A Implementation details
The source code is available at https://github.com/nkanak/brag-fake-news-campaigns.
Appendix B Experimental results
We provide the numerical results for the figures of the manuscript. Table 7 summarizes the results of Figure 2.
Table 7. Evaluation results of a framework variant with respect to different model sizes ranging from 1B to 70B. The best results for each metric are highlighted in bold.
Parameters Precision Recall F1 ROC AUC
1B 0.1022 0.9826 0.1852 0.5054
3B 0.1044 0.9751 0.1886 0.5166
8B 0.1299 0.9303 0.2280 0.6143
27B 0.1682 0.9502 0.2858 0.7106
70B 0.4238 0.8507 0.5658 0.8602
Appendix C Evaluation metrics
In this section, we expand on the main evaluation metrics used to assess the performance of our framework. The goal of these metrics is to evaluate the framework’s ability to identify coordinated disinformation campaigns effectively, particularly under extreme class imbalance.
C.1. Precision
Precision is the fraction of correctly predicted positive cases out of all cases predicted as positive by the model. In the task of disinformation detection, precision is crucial because it measures the system’s ability to minimize false positives (i.e., legitimate content incorrectly identified as part of a fake news campaign). High precision indicates that the model rarely misclassifies negative examples as positive. Mathematically, precision is defined as:
(7) $\text{Precision}=\frac{TP}{TP+FP}$
where $TP$ is the number of true positives (correctly identified disinformation campaigns) and $FP$ is the number of false positives (legitimate content wrongly classified as disinformation).
A high precision score means that when the model identifies content as part of a coordinated disinformation campaign, it is usually correct. However, high precision alone may not be sufficient, as it does not account for the model’s ability to identify all positive cases, which is measured by recall.
C.2. Recall
Recall is the fraction of true positive cases correctly identified by the model from the total actual positive cases. This metric is particularly important in disinformation detection tasks, where identifying false negatives (missed fake news campaigns) is critical. It is defined as:
(8) $Recall=\frac{TP}{TP+FN}$
where $TP$ is the number of true positives and $FN$ is the number of false negatives. A high recall indicates that the model successfully identifies a large portion of positive examples.
C.3. F1 score
The F1-score provides a harmonic mean between precision and recall, balancing the trade-off between false positives and false negatives. In the context of disinformation detection, this metric helps determine the overall accuracy of the model, accounting for both the sensitivity to true positives and the specificity to false positives. The F1-score is given by:
(9) $F1=\frac{2\times TP}{2\times TP+FP+FN}$
This score is particularly useful in our case, where the dataset suffers from class imbalance, ensuring that both precision and recall are equally weighted.
C.4. ROC AUC score
The Receiver Operating Characteristic (ROC) curve is a plot that illustrates the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings. The area under this curve (AUC) measures the overall ability of the model to distinguish between positive and negative classes. The ROC AUC score is computed as:
(10) $AUC=\int_{0}^{1}TPR(FPR)d(FPR)$
where $TPR$ is the true positive rate, and $FPR$ is the false positive rate. A higher AUC indicates a better model performance in distinguishing between classes.
C.5. False positive rate (FPR)
The false positive rate is the ratio of incorrectly predicted positive instances among all actual negative instances. In the disinformation detection task, a high FPR would mean that legitimate content is mistakenly classified as part of a coordinated campaign, which can undermine the reliability of the system. FPR is calculated as:
(11) $FPR=\frac{FP}{FP+TN}$
where $FP$ is the number of false positives and $TN$ is the number of true negatives. We aim to minimize this rate to reduce the impact of false alarms.
C.6. Class imbalance and metric interpretation
Due to the extreme class imbalance inherent in disinformation datasets, with coordinated disinformation campaigns forming a small fraction of the overall content, metrics like accuracy become less informative. Therefore, we focus on recall, F1-score, ROC AUC, and FPR, as these provide a clearer picture of the model’s performance in identifying the minority class (disinformation campaigns).
C.7. Metric performance in our framework
Our proposed framework outperforms traditional baselines, achieving high recall and F1 scores, even under significant class imbalance, as detailed in Section 3.4. The combination of network-informed prompt engineering and retrieval-augmented generation (RAG) allows our framework to maintain robust performance across all metrics without requiring additional training. The table in Section 3.4 showcases the full set of results for each metric.
Appendix D Prompt examples
In this section, we present a list of prompt examples using different prompting and RAG techniques. The list of prompting techniques includes zero-shot, few-shot and chain-of-thought prompting. These examples were retrieved during our experimentations.
D.1. Zero-shot prompting
Below is an example of zero-shot prompts used in the framework.
Please analyze the following text and its associated propagation chain that spreads on the internet like a tree-based graph.
Then, determine if it shows signs of being part of a political fake news campaign.
Respond with only "Fake" if it appears to be a political fake news campaign, or "Real" if it does not.
Text to analyze: "[TEXT]"
Tree-based graph: "(2->1), (3->2), (4->2), (5->2), (6->3), (7->2), (8->2), (9->2), (10->7), (11->2), (12->5), (13->1)"
D.2. Few-shot prompting
Here is an examples used for few-shot prompting.
You are an intelligent classifier capable of identifying political fake news campaigns. Your task is to analyze the propagation of a given tweet, focusing on aggregated network metrics to detect any coordinated behavior that might indicate a political fake news campaign. You will be provided with the tweet description and a set of statistics derived from its retweet propagation graph. These statistics summarize key structural and behavioral aspects of the network, such as the average degree centrality, eigenvector centrality, clustering coefficient, and retweet timing patterns.
The following centrality metrics are important for your analysis:
Degree centrality measures how many direct connections a user (node) has. A high degree centrality indicates that the user is directly connected to many others, suggesting they play a central role in propagating the tweet. In a coordinated campaign, several users with unusually high degree centrality may indicate organized amplification.
Eigenvector centrality measures the influence of a user within the network. It considers not only the number of connections a user has, but also the importance of those connections. A user with high eigenvector centrality is connected to other highly influential users. In coordinated behavior, high eigenvector centrality across specific users may indicate a core group driving the retweet campaign.
Retweet timing patterns refer to how the retweets are distributed over time. In coordinated campaigns, retweets often occur in rapid bursts, indicating synchronized activity. In contrast, organic propagation typically shows more dispersed retweet times.
Base your decision also on the text content, language use, and messaging tone.
Additionally, here are some labeled examples of tweet propagations to help guide your analysis. The labels indicate whether the tweet was part of a "Fake" coordinated political campaign or "Real" organic behavior.
Example 1:
Tweet text: "[TEXT]"
Propagation statistics:
Average degree centrality: 0.05
Eigenvector centrality: 0.70
Median retweet time: 12627.00 seconds
Label: Fake
Example 2:
Tweet text: "[TEXT]"
Propagation statistics:
Average degree centrality: 0.18
Eigenvector centrality: 0.58
Median retweet time: 422.00 seconds
Label: Fake
Now, analyze the following input and determine whether the tweets propagation suggests a "Fake" coordinated effort or "Real" organic behavior.
Tweet text:"[TEXT]"
Propagation statistics:
Average degree centrality: 0.10
Eigenvector centrality: 0.70
Median retweet time: 7053.00 seconds
Output:
Fake or Real.
D.3. Chain-of-thought prompting
Chain-of-thought prompting breaks down complex decisions into multiple steps, helping the model reason more effectively.
You are an intelligent classifier capable of identifying political fake news campaigns. Your task is to analyze the content of a given tweet and its associated graph metrics to detect any signs that might indicate it is part of a coordinated political fake news campaign.
Base your decision on the text content, language use, messaging tone, and the provided graph metrics. Look for signs such as sensationalism, misleading claims, highly partisan language, attempts to provoke strong emotional responses, or unusual patterns in the tweets spread as indicated by the graph metrics.
Analyze the following input tweet and its graph metrics, and determine whether it suggests a "Fake" coordinated effort or "Real" organic behavior. Please provide your reasoning step by step before giving your final decision.
Tweet text: "[TEXT]"
Graph Metrics:
- Number of nodes: 20
- Number of edges: 19
- Average degree: 1.90
- Graph density: 0.1000
- Clustering coefficient: 0.0000
Output:
Step 1: [Analyze the language and tone of the tweet]
Step 3: [Examine the graph metrics and what they suggest about the tweets spread]
Step 4: [Consider the relationship between the text content and the graph metrics]
Final Decision: [Your decision MUST be either "Fake" or "Real"]
D.4. Balanced RAG
Our proposed balanced RAG can generate pairwise examples consisting of both the positive and the negative. Here is one example of it.

### v2

@@ Abstract.
organized disinformation campaign detection; disinformation spread; fake news detection; prompt engineering; graph-aware prompt engineering; retrieval-augmented generation; class imbalance; large language models; graph classification
journalyear: 2025copyright: acmlicensedconference: Companion of the 16th ACM/SPEC International Conference on Performance Engineering; April 28-May 2, 2025; Sydney, NSW, Australia.booktitle: Companion of the 16th ACM/SPEC International Conference on Performance Engineering (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1331-6/25/04doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Discourse, dialogue and pragmatics
@@ 1. Introduction
To address the above issues, in this paper, we propose a novel framework for identifying coordinated disinformation campaigns relying only on frozen LLMs. The proposed framework encodes both the structural and textual information of a propagation tree as a text, ready to be prompted to an LLM. Using prompt engineering and retrieval-augmented generation (RAG) allows LLMs to detect organized astroturf campaigns on social networks such as $\mathbb{X}$. Our framework does not involve any training or fine-tuning of the given LLM. In contrast, it capitalizes on the power of prompt engineering and RAG techniques and enables LLMs to overcome the effects of class imbalance, thereby effectively identifying coordinated political campaigns. Our framework is flexible and can be coupled with any prompt engineering and RAG technique (e.g. few-shot prompting) (Fang et al., 2024). Furthermore, it is compatible with any decoder-only LLM.
@@
Research questions. In addition, we answer the following research questions (Section B):
@@ 3.4. Prompt engineering
Finally, a combination of prompting techniques can also be followed. We found that mixing role prompting, CoT and few-shot prompting yields the best results (Cheng et al., 2024). A list of prompt examples considering the problem of detecting astroturf campaigns is available in Appendix C. We refer to Section 2 for a more thorough explanation of prompt engineering techniques.
@@ 3.5. Balanced Retrieval-Augmented Generation (Balanced RAG)
These balanced pairs are then incorporated into the few-shot prompting approach used to query the LLM. By providing the LLM with representative examples that are both similar to the query and equally cover both classes, we enhance the model’s ability to distinguish between positive and negative instances. This approach enables the LLM to learn the critical differences that define each class, thereby improving its ability to correctly classify new, unseen propagation trees (Au et al., 2025).
@@ 4.2. Setup
Evaluation metrics. Since it is important to identify the positive samples of a dataset (i.e., fake news campaigns), we selected precision, recall, F1 score and ROC AUC scores as our main evaluation metrics.
@@ 4.5. Hyperparameter tuning
Figure 2. Comparison of different metrics (precision, recall, F1 score, and ROC AUC) across various (Llama) model sizes. The graph illustrates how each metric changes as the number of model parameters increases.
@@ 5.2. Impact of the encoded information
6. Conclusion
Appendix A Implementation details
The source code is available at https://github.com/nkanak/brag-fake-news-campaigns. Due to X’s privacy policies, the utilized dataset cannot be made publicly available. Instructions on how to download the dataset are provided in the GitHub repository.
Appendix B Discussion
@@
Appendix C Balanced RAG prompt example
Our proposed balanced RAG can generate pairwise examples consisting of both the positive and the negative.